OpenStudy (anonymous):

Someone help me with dividing cube roots--i can display it better in equation. Someone help me with dividing cube roots--i can display it better in equation. @Mathematics

OpenStudy (anonymous):

\[\sqrt[3]{x^3y^2z^7}\div \sqrt[3]{xy^4}\]

OpenStudy (asnaseer):

\[\sqrt[3]{x^3y^2z^7}\div \sqrt[3]{xy^4}=\frac{\sqrt[3]{x^3y^2z^7}}{\sqrt[3]{xy^4}}\]\[=\sqrt[3]{\frac{x^3y^2z^7}{xy^4}}=\sqrt[3]{\frac{x^2z^7}{y^2}}\]

OpenStudy (asnaseer):

this can be simplified further if you wish.

OpenStudy (anonymous):

Yes, could you try that? I found the answer and it's actually:\[z^2\sqrt[3]{x^2yz}/y\]

OpenStudy (asnaseer):

\[=\sqrt[3]{\frac{x^2z^7}{y^2}}=\sqrt[3]{\frac{x^2z^6zy}{y^3}}=\frac{z^2\sqrt[3]{x^2yz}}{y}\]

OpenStudy (asnaseer):

the steps I performed were: 1. z^7 = z * z^6 2. 1/y^2 = y/y^3 3. cube root of z^6 = z^2 4. cube root of 1/y^3 = 1/y do you follow?

OpenStudy (anonymous):

Yes, sorry i didnt reply sooner. I hope you're on here often! :)

OpenStudy (anonymous):

Thanks, do you think you can help me with domain? It might be a little hard to explain but i'll try.

OpenStudy (anonymous):

I must find the domains of f + g and f-g f(x)=-2x^(2/3) g(x)=7x^(2/3)

OpenStudy (anonymous):

So basically i have to know the domains of 5x^(2/3) and -9^(2/3) :) think you can help?